# PyraTrans: Learning Attention-Enriched Multi-Scale Pyramid Network from Pre-Trained Transformers for Effective Malicious URL Detection

### 01 Dec 2023 v1

@@ V-D Evaluation against Adversarial Attacks
As reported in the Figure 7 and TableVII, baseline methods experience notable performance degradation, whereas our model maintains stability with an AUC of 98.39%, surpassing URLNet by approximately 20 percentage points. At a fixed FPR of 0.01, our method achieves nearly 75% TPR, more than twice the TPR of baseline methods, both of which fall below 30%. These findings underscore the robustness of our method in real-world scenarios, effectively reducing false positives in security services even amid adversarial sample attacks.
@@ V-E Case Study
It is essential to evaluate a model in real-world phishing scenarios to understand its practical value. Therefore, we crawled the latest active phishing URLs reported and verified on PhishTank in November 2023 and conducted tests using models trained on the 30% Grambeddings dataset. Our model detects all malicious URLs with 100% accuracy, while Grambeddings misclassifies 7 URLs, achieving 76% accuracy, and URLNet misclassifies 4, resulting in 86% accuracy. Table VIII shows the misclassified URLs. URLNet tends to misclassify shorter URLs with fewer character features, while Grambeddings only achieves correct classification for URLs with explicitly malicious features. Grambeddings underperforms in real-world applications compared to its performance on the self-created dataset, showing a high error rate. N-grams break text into fixed-size chunks, leading to a wider and more uniform distribution of noise in the feature space, which makes it sensitive to minor fluctuations. Moreover, the Grambeddings method lacks specific feature refinement mechanisms, unlike URLNet, which uses max pooling within its CNN structure. Max pooling, though a simple operation, is effective in filtering out noise, allowing the model to focus on salient aspects of the data.
Our proposed approach uses multi-scale Learning and pyramid spatial attention mechanism, unlike the two baseline methods. This enables learning feature importance and filtering noise across every dimension of the input representation space. As a result, our method shows improved accuracy in practical scenarios, especially in detecting challenging, hard-to-distinguish, or deceptive samples. This capability indicates a significant advancement in handling complex classification tasks in real-world applications.
@@ VI Discussion
We succinctly outline our methodology through several pivotal aspects:
@@
    End-to-End Architecture: Our method employs an end-to-end three-stage approach for inference, eliminating the need for expert manual feature engineering. Firstly, a subword tokenizer extracts the tokens from the URL. Secondly, transformer-based model generates two types of embedding vectors for the unknown URL and performs further processing. Finally, a classifier predicts the result, simplifying the feature engineering process and enhancing semantic understanding.
@@
    Enhanced Learning Patterns through Hybrid Character and Word Embeddings: Our model, utilizing character and word embeddings, achieves generalization to unknown URLs and overcomes constraints observed in prior dual-path networks (e.g., URLNet), such as insensitivity to changes in URL structures and semantics, and excessive reliance on training data scale. Our dual-channel pretrained model effectively addresses these issues.
@@
    Stability and Robustness: Our method overcomes the limitations of existing models in challenging scenarios. For example, Grambeddings struggles to accurately identify different types of malicious URLs in multi-class settings, while URLNet fails to detect any malicious samples in highly imbalanced datasets with small training sets, and both models show subpar performance under adversarial attacks and cross dataset test. In contrast, our approach exhibits significant improvements in all of these scenarios.
  4. 4.
    Advanced Real-world Usability: Our model exhibits exceptional performance on the latest reported URL dataset crawled from PhishTank, correctly identifying all malicious URLs. In contrast, baseline methods perform poorly in detecting real-world dynamic URLs, particularly those with fewer features. These results demonstrate the significant practical value of our model.
  5. 5.
    Streamlined Training and Efficient Deployment: Our model achieves peak performance in only 5 training epochs—an exceptional improvement compared to the standard 20-30 epochs required by baseline methods. This remarkable efficiency is attributed to the efficient transfer of knowledge from the pretrained model to URL contexts and the integration of the lightweight multi-scale attention module. Furthermore, our model achieves optimal execution on a single GPU with over 20GB memory, delivering powerful performance while mitigating the need for high computational resources.

### 06 Dec 2023 v2

@@ V-D Evaluation against Adversarial Attacks
As illustrated in Figure 7 and Table VII, baseline methods experience a significant decline in performance under adversarial sample attacks. The accuracy of URLNet decreases to 76.10%, and that of Grambeddings to 70.18%, while our model maintains an accuracy above 90%, with an AUC of 98.39%, exceeding URLNet by about 20%. At a fixed FPR of 0.01, our method achieves a TPR of nearly 75%, more than double the TPR of baseline methods, both under 30%. These results indicate the robustness of our approach to adversarial sample attacks, suggesting increased effectiveness in preventing malicious attack evasion in real-world scenarios.
@@ V-E Case Study
We conduct a series of case studies applying our detection model to active malicious web pages to evaluate its practical utility. In November 2023, we crawled 30 active phishing URLs reported and verified on PhishTank and tested them with our model trained on 30% of the Grambeddings dataset. For comparison, tests were also conducted with the best-performing URLNet and Grambeddings models, trained on the same dataset. Results indicate that our model detected all malicious URLs with 100% accuracy, while Grambeddings misclassified 7 URLs, resulting in 76% accuracy, and URLNet misclassified 4 URLs, with an accuracy of 86%. Table VIII lists URLs that were incorrectly classified, in order to provide a detailed perspective on the performance of each model.
URLNet failed in detecting four malicious URLs, which had relatively short strings. Given that benign URLs on the internet are typically simple, URLNet might have mistakenly classified these short malicious URLs as benign due to their length similarity. This suggests that URLNet relies excessively on URL string length for classification, demonstrating limited capability in recognizing semantics and specific patterns in real-life scenarios. Conversely, Grambeddings, which combines convolutional neural networks, long short-term memory networks, and attention layers, is a more complex system. It misclassified instances of both longer and shorter URLs, indicating a minor influence of URL length. However, the significant diversity among the seven misclassified malicious URLs implies that Grambeddings’ performance could be affected by various factors, and its learning system might not have developed sufficiently generalized discriminative patterns.
In contrast, only our method demonstrated consistent or even improved performance in real-world applications, showing its adaptability to various forms of malicious URLs. Analyzing the differences at a technical level, we attribute our approach’s distinct advantage over others to its comprehensive feature consideration,includeing character-aware token embeddings, multi-level and multi-scale feature learning, and regional-level attention.
@@ VI Discussion
The proposed method, validated through a comprehensive set of experiments, has demonstrated robust, accurate, and reliable performance. Here, we briefly discuss the beneficial advantages and insights brought about by our proposed approach.
@@
    End-to-End Architecture: ur proposed method is an end-to-end network utilizing pretrained CharBERT, which requires no manual feature initialization. It directly processes raw URLs and generates character-aware subword token embeddings. In contrast, previous studies typically necessitated manual initialization of character and word-level representations and relied on dual-path neural networks. Our approach streamlines the processing workflow while maintaining efficient feature extraction capabilities.
@@
    Evaluation Metrics: Our experiments show that the proposed method significantly improves accuracy, robustness, and generalizability, consistently delivering stable and effective detection across various testing scenarios. Meanwhile, leading baseline methods, such as URLNet and Gramembedings, displayed clear weaknesses: URLNet struggled in class-imbalanced scenarios with small datasets, while Gramembedings faced significant performance fluctuation. Furthermore, these methods varied in their effectiveness in generalization and adversarial robustness tests. This highlights the need for a comprehensive performance evaluation system that goes beyond specific experimental setups.
@@
    Case Studies: Previous research often overlooked the importance of case studies, but our work emphasizes the necessity of applying models directly to active malicious links to accurately reveal their real-world performance. Our case studies have shown that even methods excelling in experimental settings can face significant challenges in practical applications. Additionally, case studies offer an opportunity to thoroughly analyze a model’s feature learning capabilities and shortcomings in information capture patterns, allowing for a more comprehensive assessment of the model’s technical design.